OpenStudy (anonymous):

What is the solution to the equation 7^ -3x ≈ 9 ?

OpenStudy (anonymous):

I'm going to assume this is 7^(-3x)=9. First, we take a log base 7 of both sides. That gives us -3x=1.129. We divide both sides by -3 and get x=-.376, our final answer.

OpenStudy (campbell_st):

I don't think you need to take the base 7 log of both sides... use base e or base 10 so you get \[\ln(7^{-3x} ) = \ln (9)\] applying the log law for powers you get \[-3x \times \ln(7) = \ln(9)\] divide both sides by ln(7) \[-3x = \frac{\ln(9)}{\ln(7)}\] then solve for x the answer above is correct... but use a common log.... not base 7

OpenStudy (anonymous):

Yeah, that would make more sense. Common logs are a lot easier to use.